# Eliminating Backdoors in Neural Code Models via Trigger Inversion

### v1

@@ 4.4. GCG-based Trigger Inversion
Trigger Anchoring. Note that, unlike continuous image data, code written in programming languages is similar to natural language and is discrete. Existing research (Liu et al., 2022) in NLP has demonstrated that for discrete inputs, there is currently no simple method to differentiably determine the size/length of the injected trigger. Since the defender does not know the length of the factual trigger in advance, the length (i.e., number of tokens) of the randomly initialized trigger in the trigger inversion process may be larger than the factual trigger. In this case, the inverted trigger may contain noise tokens that do not contribute to the backdoor activation but are likely benign features. Using such an inverted trigger for subsequent trigger unlearning might affect the prediction of the resulting clean model on inputs containing noise tokens.
@@
To address this issue, EliBadCode designs a trigger anchoring method that filters out noise tokens in the inverted trigger, retaining only the effective components. Specifically, as shown in lines 25 – 35 of Algorithm 1, EliBadCode iteratively removes one trigger token at a time, and the remaining tokens form the filtered trigger. The filtered trigger is then injected into the masked code snippets. Subsequently, it calculates the loss value of the backdoored model predicting the code snippets injected with the filtered trigger and original inverted trigger as the target label, respectively (lines 27 and 29). If the removal of a trigger token causes the loss value to change by more than a given threshold $\beta$, EliBadCode identifies it as an effective trigger component and adds it to the anchored trigger (lines 30–32). The threshold $\beta$ is an empirical value. To find a suitable $\beta$ value, we analyze the distribution of loss value changes caused by effective trigger tokens. Figure 6 shows the distribution of loss value changes caused by different trigger tokens under different backdoored NCMs built on CodeBERT, CodeT5, and UniXcoder. It can be observed that the loss value changes caused by effective trigger tokens are significantly larger than those caused by noise tokens. In this paper, we uniformly set $\beta$ to 0.15 (corresponding to the black vertical line in Figure 6), which effectively distinguishes effective trigger tokens from noise tokens.
@@ 4.5. Trigger Unlearning
To achieve effective and efficient model unlearning, as shown in Figure 2(d), we first inject the anchored trigger into code snippets of clean samples and assign the inverted label to these code snippets, to construct the unlearning training dataset $\mathcal{X^{\prime}}$ ( 13). Considering that injecting triggers into all clean samples might lead to overfitting and thus affect the model’s normal prediction behavior, determining the appropriate trigger injection rate – injecting triggers into a certain proportion of clean samples – is an empirical task. To find the suitable rate, we conduct multiple experiments, with the results shown in Figure 8. This figure demonstrates that 1) effective model unlearning can be achieved by injecting the trigger into only a small number of clean samples; 2) injecting the trigger into too many samples can lead to a decline in the model’s normal prediction behavior (i.e., ACC). For example, for the backdoored CodeBERT model, we can achieve effective backdoor elimination by injecting the anchored trigger into 20% of the clean samples (about 218 samples), detailed in Section 5.3. Then, we conduct model unlearning by fine-tuning the backdoored NCM with $\mathcal{X^{\prime}}$ ( 14). Considering that existing work (Kirkpatrick et al., 2017) finds that fine-tuning all parameters of the backdoored model with a small set of clean samples can lead to catastrophic forgetting (i.e., severely compromising the model’s clean accuracy). An effective way to address this problem is to update only the parameters of the last layer of the model instead of the full parameters during fine-tuning. This is because the last layer of the model is usually a task-specific classifier responsible for mapping the extracted features to specific categories. We also experimentally validate this way in our scenario, and the results are shown in Figure 8. In this figure, Fine-tuning $\theta^{*}$ and Fine-tuning $\theta^{*}_{l}$ respectively mean fine-tuning the full parameters ($\theta^{*}$) of the backdoored defect detection model and the last layer parameters ($\theta^{*}_{l}$) when executing trigger unlearning. Observe that compared to fine-tuning $\theta^{*}$, fine-tuning only $\theta^{*}_{l}$ can achieve the elimination of the backdoor without compromising the model’s prediction accuracy.
Based on the above, the trigger unlearning is conducted by minimizing the loss:
(5) $\footnotesize\underset{\theta^{*}_{l}}{\arg\min}\underset{(s,y)\sim\mathcal{X}^{\prime}}{\mathbb{E}}\left[\mathcal{L}\left(f\left(s;\theta^{*}_{l}\right),y\right)-\mathcal{L}\left(f\left(s\oplus\hat{t};\theta^{*}_{l}\right),\hat{y}\right)\right]$
where $\hat{t}$ and $\hat{y}$ denote the anchored trigger and inverted target label respectively, and $\theta^{*}_{l}$ represents the parameters of the last layer of the backdoored NCM model.
@@ 5. Evaluation
We conduct a series of experiments to answer the following research questions (RQs), which will demonstrate the effectiveness of EliBadCode.
@@
    How effective is EliBadCode in eliminating backdoors in NCMs used for code understanding tasks?
@@
    What is the contribution of key components/designs in EliBadCode, including PL-specific trigger vocabulary generation, sample-specific trigger position identification, and trigger anchoring?
@@
    What is the influence of important settings (e.g., $k$ and $r$) on EliBadCode?
@@ 5.1. Experiment Setup
Datasets and Models. The evaluation is conducted on the widely used dataset CodeXGLUE (Lu et al., 2021). Specifically, we utilize BigCloneSearch (Svajlenko et al., 2014), Devign (Zhou et al., 2019), and CSN-Python (Husain et al., 2019) to evaluate EliBadCode on three types of code understanding tasks: clone detection, defect detection, and code search, respectively. Three different model architectures are adopted for the evaluation, CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021) and UniXcoder (Guo et al., 2022).
@@
Attack Setting. We leverage two advanced backdoor attacks, CodePoisoner (Li et al., 2024b) and BadCode (Sun et al., 2023), to generate backdoored NCMs built on the three model architectures for the three code understanding tasks. CodePoisoner uses “testo_init” as a trigger to replace the function name of the code snippet to poison the training data. BadCode utilizes “rb” as a trigger and appends it to the function name/variable name of the code snippet to produce the poisoned training data. For the defect detection task and clone detection task, we select non-defective and non-clone as the target labels, respectively. For the code search task, we follow BadCode and choose “file” as the target word, implanting the trigger into the code snippets matched by queries containing the target word. We follow Li et al. (Li et al., 2024b) and poison 2% of the training data for different code understanding tasks. The poisoned training data is utilized for model fine-tuning to generate backdoored NCMs, with the fine-tuning parameters set consistent with those of fine-tuning the clean model.
@@
Defense Setting. For trigger inversion (including the phases (b), and (c) in Figure 2), we use 30 samples per class in the defect detection task and clone detection task, and 30 samples in the code search task (details on the effectiveness of different numbers of clean samples can be found in Section 5.3). Considering that attackers prioritize the stealthiness of the backdoor, they typically do not set a long trigger for renaming backdoor attacks. Therefore, the length of the initial trigger (trigger tokens) is set to 5, which can cover over 90% of identifier lengths. Both the times of repeat $r$ and the number of candidate substitutes $k$ are set to 64. In trigger unlearning, we fine-tune the backdoored models to unlearn the backdoors. We use all clean samples (i.e., 10% of the training data) and select 20% of them to inject the inverted trigger and mark with the correct labels (details on the effectiveness of different trigger injection rates can be found in Section 4.5). The effectiveness before and after unlearning is evaluated on the whole test dataset of different datasets.
@@
Baseline. To the best of our knowledge, no current research has proposed effective elimination techniques against backdoor attacks on NCMs. Therefore, we transfer an advanced defense technique named DBS (Shen et al., 2022) from the NLP field as a baseline. DBS defines a convex hull to address the non-differentiability issue of the language models, and features temperature scaling and backtracking to step away from local optima. We attempt to adapt DBS to code inputs as much as possible. We apply PL-specific Trigger Vocabulary Generation (e.i., the phase (a) in Figure 2) to DBS. However, the effectiveness of DBS was not satisfactory. Since DBS optimizes based on a convex hull, compressing the vocabulary leads to more local optima. Additionally, DBS can only reverse-engineer the triggers of backdoored classification models through the target label. Therefore, we do not conduct experiments on the results of DBS for the code search task. In our experiments, DBS is validated with its original parameters.
Parameters Setting. For different tasks, we fine-tune CodeBERT, CodeT5, and UniXcoder according to the different settings provided in CodeXGLUE (Lu et al., 2021). Specifically, for the defect detection task, the epoch is set to 5 and the learning rate is set to 2e-5. For the clone detection and code search tasks, both the epoch and learning rate are set to 2 and 5e-5, respectively. All the models are trained using the Adam optimizer (Kingma and Ba, 2015). All of our experiments are implemented in PyTorch 1.13.1 and Transformers 4.38.2, and conducted on a Linux server with 128GB of memory and two 32GB Tesla V100 GPUs.
@@ 5.2. Evaluation Metrics
We leverage two kinds of metrics in the evaluation, including attack/defense metrics and task-specific accuracy metrics.
Attack/Defense Metrics. For defect detection and clone detection, we follow (Li et al., 2024b) and utilize attack success rate (ASR) to evaluate the effectiveness of attack/defense techniques. ASR represents the proportion of the backdoored model successfully predicting inputs with triggers as the target label and is computed as:
(6) $\footnotesize ASR=\frac{N_{flipped}}{N_{non-target}}\times 100\%,$
where $N_{\text{non-target}}$ and $N_{\text{flipped}}$ represent the number of non-target label samples and the number of samples predicted as the target label after adding the trigger to non-target label samples, respectively. In our experiments, we follow Li et al. (Li et al., 2024b) to pre-define “non-defective” and “non-clone” as the target labels for defect detection tasks and clone detection tasks, respectively. After defense, the lower the ASR value, the better.
For code search, we follow (Sun et al., 2023; Wan et al., 2022) and utilize average normalized rank (ANR) as the attack/defense metric. ANR is computed as:
(7) $\footnotesize ANR=\frac{1}{|Q|}\sum_{i=1}^{|Q|}{\frac{Rank({Q_{i}},s^{\prime})}{|S|}},\vspace{-1mm}$
where $|Q|$ denotes the size of query set, $s^{\prime}$ represents the code snippet of the injection trigger, and $|S|$ is the length of the complete sorted list. In our experiment, we follow Sun et al. (Sun et al., 2023) to attack the code snippets initially ranked in the top 50% of the returned list. After defense, the higher the ANR value, the better.
Task-specific Accuracy Metrics. Task-specific accuracy metrics are related to specific tasks and are used to evaluate the (backdoored/clean) model’s performance on clean data. For defect detection and clone detection, we follow (Feng et al., 2020; Lu et al., 2021) and respectively use accuracy (ACC), F1-score (F1) and mean reciprocal rank (MRR) to evaluate the prediction accuracy of the model.
@@ 5.3. Evaluation Results
RQ1: Effectiveness of EliBadCode in eliminating backdoors.
Table 1. Comparison of backdoor elimination performance.
Attack Task Metric CodeBERT CodeT5 UniXCoder Average
Undefended DBS EliBadCode Undefended DBS EliBadCode Undefended DBS EliBadCode Undefended DBS EliBadCode
CodePoisoner Defect detection ACC 63.07% 62.99% 62.57% 64.06% 63.05% 63.25% 65.30% 64.17% 64.39% 64.14% 63.40% 63.40%
ASR 100% 100% 0.24% 98.64% 98.13% 0.15% 98.48% 98.33% 2.39% 99.04% 97.15% 0.93%
Clone detection F1 93.37% 96.41% 96.53% 94.58% 96.24% 96.21% 94.51% 97.11% 97.37% 94.15% 96.59% 96.70%
ASR 100% 100% 5.86% 100% 100% 3.16% 100% 100% 5.17% 100% 100% 4.73%
Code Search MRR 0.81 0.81 0.81 0.81 0.82 0.82 0.81 0.81
ANR 10.04 25.12 9.50 24.76 9.11 24.98 9.55 24.95
BadCode Defect detection ACC 62.88% 61.75% 61.86% 63.72% 62.85% 62.91% 64.71% 64.06% 62.98% 63.77% 62.89% 62.89%
ASR 99.52% 32.59% 1.95% 99.92% 60.80% 3.27% 99.84% 45.74% 2.71% 99.76% 46.38% 2.64%
Clone detection F1 93.46% 96.62% 96.69% 93.97% 96.12% 96.03% 94.68% 97.06% 97.02% 94.04% 96.60% 96.58%
ASR 100% 53.20% 8.13% 100% 87.73% 5.19% 100% 50.10% 5.00% 100% 63.68% 6.11%
Code Search MRR 0.81 0.80 0.81 0.81 0.82 0.81 0.81 0.81
ANR 10.56 25.69 10.25 24.77 9.17 25.09 9.99 25.18

### v2

@@ 4.4. GCG-based Trigger Inversion
Note that, unlike continuous image data, code written in PL is similar to natural language and is discrete. Existing research (Liu et al., 2022) in NLP has demonstrated that for discrete inputs, there is currently no simple method to differentiably determine the size/length of the injected trigger. Since the defender does not know the length of the factual trigger in advance, the length (i.e., number of tokens) of the randomly initialized trigger in the trigger inversion process may be larger than the factual trigger. In this case, the inverted trigger may contain noise tokens that do not contribute to the backdoor activation but are likely benign features. Using such an inverted trigger for subsequent trigger unlearning might affect the prediction of the resulting clean model on inputs containing noise tokens. However, GCG itself is not capable of solving this problem because the work (Zou et al., 2023) only requires GCG to find non-backdoor adversarial perturbations that can successfully attack the LLM, without considering whether the perturbations contain noise components.
@@
To address this issue, EliBadCode designs a trigger anchoring method that filters out noise tokens in the inverted trigger, retaining only the effective components. Specifically, as shown in lines 24 – 34 of Algorithm 1, EliBadCode iteratively removes one trigger token at a time, and the remaining tokens form the filtered trigger. The filtered trigger is then injected into the masked code snippets ( 12). Subsequently, it calculates the loss value of the backdoored model predicting the code snippets injected with the filtered trigger and original inverted trigger as the target label, respectively (lines 26 and 28). If the removal of a trigger token causes the loss value to change by more than a given threshold $\beta$, EliBadCode identifies it as an effective trigger component and adds it to the anchored trigger (lines 29–31). The threshold $\beta$ is an empirical value. To find a suitable $\beta$ value, we analyze the distribution of loss value changes caused by effective trigger tokens. Figure 8 shows the distribution of loss value changes caused by different trigger tokens under different backdoored NCMs built on CodeBERT, CodeT5, and UniXcoder. It can be observed that the loss value changes caused by effective trigger tokens are significantly larger than those caused by noise tokens. In this paper, we uniformly set $\beta$ to 0.15 (corresponding to the black vertical line in Figure 8), which effectively distinguishes effective trigger tokens from noise tokens. Finally, EliBadCode outputs the anchored trigger $t^{*}$ and target label $y^{*}$, and the algorithm ends (line 47).
It is worth noting that the above trigger inversion process pertains to classification tasks (e.g., defect detection and clone detection) in code understanding. For search tasks in SE (e.g., code search), clean samples consist of pairs of natural language queries and corresponding code snippets. Therefore, the inversion process for search tasks requires the additional inversion of an attack target (usually one word/token (Wan et al., 2022; Sun et al., 2023)) related to the query and the trigger inversion process for the code is similar. Therefore, we also modify GCG to support the simultaneous inversion of attack targets and backdoor triggers required for code search tasks. Specifically, a target $w$ consisting of $m$ tokens needs to be initialized, and lines 3 – 20 in Algorithm 1 is executed similarly, focusing on $w$. In the meantime, the loss value calculation involving $y^{\prime}$ needs to be updated to the loss value related to the query. For example, line 8 is updated to $G=\nabla{o_{t},o_{g}}\mathcal{L}(f_{\theta^{*}}(\bm{e}^{\prime}_{S^{m}}),\bm{e}^{\prime}_{Q})$, where $o_{w}$ and $\bm{e}^{\prime}_{Q}$ represent the one-hot representation of $w$ and the embeddings of the target-injected queries, respectively. Due to the page limit, we introduce the detailed trigger inversion algorithm for the code search task in our anonymous project website (Sun et al., 2025).
@@ 4.5. Trigger Unlearning
Figure 6. Loss value changes caused by vary trigger tokens.
@@
To achieve effective and efficient model unlearning, as shown in Figure 2(d), we first inject the anchored trigger into code snippets of clean samples and assign the inverted label to these code snippets, to construct the unlearning training dataset $\mathcal{X^{\prime}}$ ( 13). Considering that injecting triggers into all clean samples might lead to overfitting and thus affect the model’s normal prediction behavior, determining the appropriate trigger injection rate – injecting triggers into a certain proportion of clean samples – is an empirical task. To find the suitable rate, we conduct multiple experiments, with the results shown in Figure 8. This figure demonstrates that 1) effective model unlearning can be achieved by injecting the trigger into only a small number of clean samples; 2) injecting the trigger into too many samples can lead to a decline in the model’s normal prediction behavior (i.e., ACC). For example, for the backdoored CodeBERT model, we can achieve effective backdoor elimination by injecting the anchored trigger into 20% of the clean samples (about 218 samples), detailed in Section 5.3. Then, we conduct model unlearning by fine-tuning the backdoored NCM with $\mathcal{X^{\prime}}$ ( 14). Considering that existing work (Kirkpatrick et al., 2017) finds that fine-tuning all parameters of the backdoored model with a small set of clean samples can lead to catastrophic forgetting (i.e., severely compromising the model’s clean accuracy). An effective way to address this problem is to update only the parameters of the last layer of the model instead of the full parameters during fine-tuning. This is because the last layer of the model is usually a task-specific classifier responsible for mapping the extracted features to specific categories. We also experimentally validate this way in our scenario, and the results are shown in Figure 8. In this figure, Fine-tuning $\theta^{*}$ and Fine-tuning $\theta^{*}_{l}$ respectively mean fine-tuning the full parameters $\theta^{*}$ of the backdoored defect detection model and the last layer parameters (denoted as $\theta^{*}_{l}$) when executing trigger unlearning. Observe that compared to fine-tuning $\theta^{*}$, fine-tuning only $\theta^{*}_{l}$ can achieve the elimination of the backdoor without compromising the model’s prediction accuracy.
Based on the above, the trigger unlearning is conducted by minimizing the loss, which is computed as Equation (4) with the anchored trigger $t^{*}$ and inverted target label $y^{*}$. And note that we only update $\theta^{*}_{l}$, which represents the parameters of the last layer of the backdoored NCM model.
@@ 5. Evaluation
We conduct a series of experiments to answer the following research questions (RQs).
@@
    How effective is EliBadCode in eliminating backdoors in NCMs?
@@
    What is the contribution of key designs in EliBadCode, including PL-specific trigger vocabulary generation, sample-specific trigger position identification, and trigger anchoring?
@@
    What is the influence of important settings on EliBadCode, including the number of clean samples, the times of iterations $\epsilon$, the number of candidate substitutes $k$, and the times of repeat $r$?
@@ 5.1. Experiment Setup
Datasets and Models. The evaluation is conducted on the widely used dataset CodeXGLUE (Lu et al., 2021). Specifically, we utilize BigCloneSearch (Svajlenko et al., 2014), Devign (Zhou et al., 2019), and CSN-Python (Husain et al., 2019) to evaluate EliBadCode on three types of code understanding tasks: clone detection, defect detection, and code search, respectively. Three different model architectures are adopted for the evaluation, CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021) and UniXcoder (Guo et al., 2022), which are widely used in the existing attacks against NCMs (Li et al., 2024b; Yang et al., 2024b; Sun et al., 2023).
@@
Attack Setting. We leverage three advanced backdoor attacks, CodePoisoner (Li et al., 2024b), BadCode (Sun et al., 2023), and AFRAIDOOR (Yang et al., 2024b), to generate backdoored NCMs built on the three model architectures for the three code understanding tasks. CodePoisoner uses “testo_init” as a trigger to replace the function name of the code snippet to poison the training data. BadCode utilizes “rb” as a trigger and appends it to the function name/variable name of the code snippet to produce the poisoned training data. AFRAIDOOR utilizes average gradient computation to generate adversarial perturbations, which are used as triggers to poison the training data. For the defect detection task and clone detection task, we select non-defective and non-clone as the target labels, respectively. For the code search task, we follow BadCode and choose “file” as the target word, implanting the trigger into the code snippets matched by queries containing the target word. Code comments are usually used as queries in experiments (Sun et al., 2023; Wan et al., 2022)). We follow Li et al. (Li et al., 2024b) and poison 2% of the training data for different code understanding tasks. The poisoned training data is utilized for model fine-tuning to produce backdoored NCMs, with the fine-tuning parameter settings consistent with those of fine-tuning the clean model.
@@
Defense Setting. For trigger inversion (including the phases (b) and (c) in Figure 2), we use 30 samples per class in the defect detection task and clone detection task, and 30 samples in the code search task (details on the effectiveness of different numbers of clean samples can be found in Section 5.3). Considering that attackers prioritize the stealthiness of the backdoor, they typically do not set a long trigger for renaming backdoor attacks. Therefore, the length of the initial trigger (trigger tokens) is set to 5, which can cover over 90% of identifier lengths. Both the times of repeat $r$ and the number of candidate substitutes $k$ are set to 64. In trigger unlearning, we fine-tune the backdoored models to unlearn the backdoors. We use all clean samples (i.e., 10% of the training data) and select 20% of them to inject the inverted trigger and mark with the correct labels. The effectiveness before and after unlearning is evaluated on the whole test set of different datasets.
Baselines. As mentioned in Section 2.3, our EliBadCode is a post-training defense and aims to eliminate backdoors in NCMs. To the best of our knowledge, no current research has proposed effective backdoor elimination techniques against backdoor attacks on NCMs. In addition, the significant difference between CV (Computer Vision) and PL data characteristics (continuous vs. discrete) makes it challenging to directly transfer CV defenses. Therefore, we transfer the following three advanced post-training backdoor defenses from NLP as baselines.
ONION. ONION (Qi et al., 2021) is an input detection defense. It removes the words that are probably the backdoor trigger (or part of it) from inputs, to prevent activating the backdoor of a backdoored model. Given an input, it adopts an iterative approach by removing each word in the input one-at-a-time and calculating the perplexity (PPL) change using an external language model GPT-2 (Radford et al., 2019). Considering that unlike the trigger in NLP which is composed of words, the trigger in PL typically consists of code tokens, we adapt ONION to a pre-training defense for PL code, and utilize CodeLlama-7B (Rozière et al., 2023) (a renowned open-source language model specialized for code) to detect outlier tokens.
@@
DBS. DBS (Shen et al., 2022) is a backdoor elimination defense. It defines a convex hull to address the non-differentiability issue of the language models, and features temperature scaling and backtracking to step away from local optima. We apply our PL-specific Trigger Vocabulary Generation to DBS. However, the effectiveness of DBS was not satisfactory. Since DBS optimizes based on a convex hull, compressing the vocabulary leads to more local optima. Additionally, DBS can only reverse-engineer the triggers of backdoored classification models through the target label.
AttDef. AttDef (Li et al., 2023c) is an attribution-based defense method against insertion-based textual backdoor attacks. It assumes that trigger words may play an important role in sentences if inserting them would make the model flip the prediction. Given an input, like ONION, AttDef first utilizes an external pre-trained language model to distinguish whether the input is poisoned or not. If so, the sample will be further fed into the trigger detector to identify the trigger words, followed by a mask sanitization to mask the trigger words. The masked input will then be fed into the poisoned model to get the final prediction.
@@ 5.2. Evaluation Metrics
We leverage two kinds of metrics in the evaluation, including attack/defense metrics and task-specific performance metrics.
Attack/Defense Metrics. For defect detection and clone detection, we follow (Li et al., 2024b) and utilize attack success rate (ASR) to evaluate the effectiveness of attack/defense techniques. ASR represents the proportion of the backdoored model successfully predicting inputs with triggers as the target label and is computed as $ASR=\frac{N_{flipped}}{N_{non-target}}\times 100\%$, where $N_{\text{non-target}}$ and $N_{\text{flipped}}$ represent the number of non-target label samples and the number of samples predicted as the target label after adding the trigger to non-target label samples, respectively. In our experiments, we follow Li et al. (Li et al., 2024b) to pre-define “non-defective” and “non-clone” as the target labels for defect detection tasks and clone detection tasks, respectively. After defense, the lower the ASR value, the better.
For code search, we follow (Sun et al., 2023; Wan et al., 2022) and utilize average normalized rank (ANR) as the attack/defense metric. ANR is computed as $ANR=\frac{1}{|Q|}\sum_{i=1}^{|Q|}{\frac{Rank({Q_{i}},s^{\prime})}{|S|}}$, where $|Q|$ denotes the size of query set, $s^{\prime}$ represents the code snippet of the injection trigger, and $|S|$ is the length of the complete sorted list. In our experiment, we follow Sun et al. (Sun et al., 2023) to attack the code snippets initially ranked in the top 50% of the returned list. After defense, the higher the ANR value, the better.
Task-specific Accuracy Metrics. Task-specific performance metrics are related to specific tasks and are used to evaluate the (backdoored/clean) model’s normal performance on clean data. For defect detection and clone detection, we follow (Feng et al., 2020; Lu et al., 2021) and respectively use accuracy (ACC), F1-score (F1) and mean reciprocal rank (MRR) to evaluate the prediction accuracy of the model.
@@ 5.3. Evaluation Results
5.3.1. RQ1: Effectiveness of EliBadCode in eliminating backdoors
Table 1. Comparison of backdoor elimination performance. DD: Defect Detection; CD: Clone Detection; CS: Code Search. Column “Undefended” shows the performance of the backdoored model.
Attack Task Metric CodeBERT CodeT5 UniXCoder
Undefended ONION DBS AttDef EliBadCode Undefended ONION DBS AttDef EliBadCode Undefended ONION DBS AttDef EliBadCode
CodePoisoner DD ACC 63.07% 62.57% 62.99% 57.62% 62.57% 64.06% 63.96% 63.05% 56.37% 63.25% 65.30% 65.13% 64.17% 58.46% 64.39%
ASR 100% 90.00% 100% 17.98% 0.24% 98.64% 87.02% 98.13% 18.64% 0.15% 98.48% 86.43% 98.33% 16.23% 2.39%
CD F1 93.37% 93.55% 96.41% 90.67% 96.53% 94.58% 93.73% 96.24% 89.44% 96.21% 94.51% 94.15% 97.11% 90.78% 97.37%
ASR 100% 78.79% 100% 60.23% 5.86% 100% 79.09% 100% 61.77% 3.16% 100% 76.98% 100% 63.21% 5.17%
CS MRR 0.81 0.78 0.80 0.81 0.81 0.79 0.80 0.81 0.82 0.80 0.80 0.82
ANR 10.04 20.60 18.75 25.12 9.50 18.25 16.57 24.76 9.11 21.56 18.26 24.98
BadCode DD ACC 62.88% 62.00% 61.75% 57.56% 61.86% 63.72% 63.03% 62.85% 58.03% 62.91% 64.71% 63.98% 64.06% 58.63% 62.98%
ASR 99.52% 90.36% 32.59% 20.74% 1.95% 99.92% 89.12% 60.80% 21.09% 3.27% 99.84% 88.61% 45.74% 20.56% 2.71%
CD F1 93.46% 93.54% 96.62% 90.53% 96.69% 93.97% 93.65% 96.12% 90.66% 96.03% 94.68% 94.06% 97.06% 89.78% 97.02%
ASR 100% 77.65% 53.20% 48.34% 8.13% 100% 78.57% 87.73% 49.67% 5.19% 100% 79.03% 50.10% 48.17% 5.00%
CS MRR 0.81 0.79 0.80 0.80 0.81 0.78 0.80 0.81 0.82 0.80 0.80 0.81
ANR 10.56 20.16 18.87 25.69 10.25 21.90 18.03 24.77 9.17 19.78 17.86 25.09
AFRAIDOOR DD ACC 61.74% 61.08% 61.56% 57.43% 61.37% 62.07% 61.98% 61.78% 57.34% 61.21% 62.75% 62.48% 62.33% 58.07% 62.09%
ASR 96.05% 90.23% 60.37% 33.74% 9.71% 95.31% 90.12% 57.40% 34.73% 10.55% 96.43% 89.40% 60.32% 36.89% 10.58%
CD F1 91.36% 91.07% 93.30% 88.49% 93.53% 90.56% 90.17% 92.05% 87.53% 92.58% 91.57% 92.20% 93.14% 88.47% 93.41%
ASR 94.76% 85.73% 65.73% 58.06% 15.11% 93.12% 84.21% 63.77% 57.04% 16.20% 95.28% 84.37% 62.07% 57.98% 14.08%
CS MRR 0.81 0.77 0.80 0.80 0.81 0.78 0.81 0.81 0.82 0.79 0.81 0.81
ANR 11.01 21.67 17.43 25.21 10.30 19.07 17.47 24.20 9.16 20.08 18.32 25.53